CAN ONE DETERMINE THE ELECTRIC DIPOLAR MOMENT OF KrH^{+} X ^{1}Σ^{+} FROM ONLY FRE-QUENCY DATA OF PURE ROTATIONAL AND VIBRATION-ROTATIONAL SPECTRA?

\underline{J. F. OGILVIE}, Universidad de Costa Rica, Escuela de Quimica, Ciudad Universitaria Rodrigo Facio, San Pedro de Montes de Oca, San Jose 2060, Costa Rica; also, Simon Fraser University, Department of Mathematics and Centre for Experimental and Constructive Mathematics, 8888 University Drive, Burnaby, British Columbia V5A 1S6, Canada.

In agreement with Molski’s derivation^{a} of an insignificant value t_{0}^{Kr} = −0.0146 ± 0.0335 of a parameter related to the rotational g factor, our objective fit of 352 appraised measurements of frequencies of pure rotational transitions and wave numbers of vibration-rotational transitions of KrH^{+} in various isotopic species ^{78,80,82,83,84,86}Kr^{1,2}H^{+} indicates that one might estimate^{b} a value of the rotational g factor if adiabatic corrections be negligible, but that no corresponding estimate of the permanent electric dipolar moment is practicable. The application of computational spectrometry^{c} to the prediction of an accurate wave number of the pure fundamental vibrational transition of LiH is also discussed.

aM. Molski, Molecular Physics, 100, 3545 - 3552 (2002)

bJ. F. Ogilvie and S. C. Liao, Chemical Physics Letters, 226, 281 - 288 (1994)

cJ. F. Ogilvie and J. Oddershede, Advances in Quantum Chemistry, 48, 253 - 317 (2005)